Problem 18-10 On March 1, 2017, Culver Construction Company contracted to construct a factory building for Fabrik Manufacturi2018 2019 (b) Using the completed-contract method, prepare schedules to compute the profit or loss to be recognized as a resu2017 2018 2019 LINK TO TEXT Question Attempts: 0 of 3 used SAVE FOR LATER SUBMIT ANSWER
Answer #1

Answer :

Given data is :

Total contract price = $ 8,450,000

a)From the given data :

Particulars 2017 2018 2019
Cost in present year 3,052,800 2,254,300 1,962,900
Total cost 3,052,800

= 3,052,800 + 2,254,300

= 5,307,100

= 5,307,100 + 1,962,900

= 7,270,000

Estimated cost 3,307,200 1,962,900 0
Total estimated cost 6,360,000

= 5,307,100 + 1,962,900

= 7,270,000

7,270,000
Percentage of completion = (cost incurred till date(Total cost) / total estimated cost)

= 3,052,800 / 6,360,000

= 0.48

= 48%

= 5,307,100 / 7,270,000

= 0.73

= 73%

= 7,270,000 / 7,270,000

= 1

= 100%

Total revenue $ 8,450,000 $ 8,450,000 $ 8,450,000
Recognized revenue total

=8,450,000 * 48%

= 4,056,000

= 8,450,000 * 73%

= 6,168,500

= 8,450,000 * 1

= 8,450,000

Recognized revenue in past year 0 4,056,000 6,168,500
Recognized revenue in present year 4,056,000

= 6,168,500 - 4,056,000

= 2,112,500

2,281,500

b)Using completed contract method :

Particulars 2017 2018 2019
Revenue in present year 4,056,000 2,112,500 2,281,500
Total cost in present year 3,052,800 2,254,300 1,962,900
Gross Profit

= 4,056,000 - 3,052,800

= 1,003,200

= 2,112,50- 2,254,300

= - 141,800

= 2,281,500 - 1,962,900

= 318,600